# NL in the Middle: Code Translation with LLMs and Intermediate Representations

### v1

@@ V Discussion
V-A Analysis of Target Language Performance
We found that code translation to Python performs worse than to C, C++, and Go. After reviewing the generated Python files with errors, we discovered that a key cause was missing critical Python imports. For example, the generated code uses ‘bufio’ or ‘bisect’, but this package is missing from the initial import statement. Hallucinations also occur where the generated code attempts to use a variable that was not defined or initialized, causing an error, or incorrectly uses a function resulting in runtime errors such as “ValueError: invalid literal for int() with base 10”. Interestingly, StarCoder has fewer of these problems and has generally better performance for Python compared to Open Gpt4 8X7B and CodeGen. A potential avenue for this insight is to leverage specific models for code translations to a specific language. To elaborate, from these results, it might be beneficial to use Open Gpt4 8X7B for translating to Go but to use StarCoder for code translation to Python.
V-B Limitations
One limitation of our work is the focus on open-source LLMs. As shown in [3], the proprietary GPT-4 model achieved significantly better results compared to the other approaches. However, we find our work to be more reflective of enterprise use cases for code translation as many companies have policies against using proprietary LLMs with sensitive data (such as company code) [44, 45, 46]. Our study shows that leveraging an open-source LLM could be beneficial for code translation and these open-source models could also operate locally, assuming the hardware requirements are met.
@@ V-B Limitations
V-C Future Work
Possible future work could be to extend these results to more code translation pairs, other code languages, and different open-source models. More code translation pairs could be obtained by incorporating the full CodeNet dataset instead of the selected sample (which we used to enable a comparison to previous work). In addition, experimentation with few-shot learning and variations of CoT prompting (e.g., Structured CoTs [25], COTTON [26]) could also be conducted to determine if intermediate representations would also aid in increasing performance for code translation. Other datasets such as CodeTransOcean and CodeTrans from CodeXGLUE could be incorporated if we write unit tests for them to check correctness of the translated code. Other languages could be investigated by asking for different target languages with the current datasets used in this paper; the testing pipeline uses generic input and output files that could be leveraged for other target languages. Using the full CodeNet dataset would also help with trying other code languages as the full CodeNet dataset also contains the source code for other languages beyond Java, C, C++, Go, and Python. Future work could also involve using CodeGeeX and checking the extensibility of our findings to different open-source models.
Other future work could be to improve the post-processing of the LLM output along with trying more specific intermediate code representations during code translation. In this work, we post-processed the LLM output to reduce the frequency of noisy non-code lines. However, we only addressed the most common issues using a simple heuristic check, but there could be more specific post-processing to boost the percentage of successful translations. Another lens of future work would be to leverage more specific intermediate code representations. In our prompt formats, we focused on NL and a general AST but there has been progress on self-optimizing ASTs to incorporate specific language nuances [47].
VI Conclusion
@@
Appendix A Summary of Programming Languages for Code Translation Benchmarks
Table IV summarizes the popular code translation benchmark datasets, with the dataset size derived from the reported value in the corresponding paper. Notably, some values are reported on a program-level whereas others are reported on a sample level. Given their different organizational structures, we omit their conversion to the reader and present the numbers simply for reference rather than comparison.
TABLE IV: Popular code translation benchmarks sorted alphabetically by name where PL = programming languages.
Dataset Name # of PL Size
AVATAR [40] 2 9.5K
CodeNet [35] 55 13.9M
CodeTrans (from CodeXGLUE [41]) 2 11.8K
Commons CLI (from [3, 48]) 2 22
Click (from [3, 49]) 2 15
CoST [50] 7 132.0K
DLTrans (from CodeTransOcean [37]) 4 408
EvalPlus (from [3, 51]) 2 164
LLMTrans (from CodeTransOcean [37]) 8 350
MultilingualTrans (from CodeTransOcean [37]) 8 30.4K
NicheTrans (from CodeTransOcean [37]) 45 236.5K
Phrase-Based Dataset [52] 2 20.5k
xCodeEval [53] 11 5.5M
XLCoST [54] 7 1.0M
Appendix B Prompts Used for Experiments
The following sections present some of the prompts used in the initial experiments (Phase 1) to test the effect of various prompt formats.
B-A Zero-Shot Prompt (0SP) [3]
[…]
2
3message = [{"role": "system", "content": "You are a helpful assistant."},
4 {"role": "user", "content": content}]
B-B One-Shot Prompt (1SP)
[…]
2
3sample_input = """
4// Example: Translate the above Python code to C.
5N, K = map(int, input().split())
6if N % K == 0:
7 print(0)
8else:
9 print(1)
10"""
11
12sample_output = """
13// Response:
14int main() {
15 int N, K;
16 scanf("%d %d", &N, &K);
17 if (N % K == 0) {
[…]
19 } else {
[…]
21 }
22 return 0;
23}
24
25// End of Code
26"""
27
28message = [{"role": "system", "content": "You are a helpful assistant."},
[…]
B-C Two-Step Prompt with an Intermediate Natural Language Summary (2S-NL)
1# Prompt 1 (output = summary)
[…]
3
4message = [
5 {"role": "system", "content": "You are a helpful assistant."},
6 {"role": "user", "content": content}]
7
8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
9
10# Prompt 2
[…]
12
13message = [{"role": "system", "content": "You are a helpful assistant."},
14 {"role": "user", "content": content}]
B-D Two-Step Prompt with AST Followed by NL (2S-AST-NL)
1 # Prompt 1 (output = summary)
[…]
3
4message = [
5 {"role": "system", "content": "You are a helpful assistant."},
6 {"role": "user", "content": content}]
7
8intermediate = llm.create_chat_completion(message) [’choices’][0][’message’][’content’]
9
10# Prompt 2
[…]
12
13message = [{"role": "system", "content": "You are a helpful assistant."},
14 {"role": "user", "content": content}]
B-E Chain-of-Thought Prompt with an Intermediate Natural Language Summary (CoT-NL)
[…]
2
[…]
4
[…]
6
7message = [{"role": "system", "content": "You are a helpful assistant."},
[…]
B-F Chain-of-Thought Prompt with AST Followed by NL (CoT-AST-NL)
[…]
2
[…]
4
[…]
6
7message = [{"role": "system", "content": "You are a helpful assistant."},
[…]

### v2

@@ V Discussion
V-A Limitations
One limitation of our work is the focus on open-source LLMs. According to Pan et al. [2], the proprietary GPT-4 model achieved significantly better results compared to the other approaches. However, we find our work to be more reflective of enterprise use cases for code translation as many companies have policies against using proprietary LLMs with sensitive data (such as company code) [42, 43, 44]. Our study shows that leveraging an open-source LLM could be beneficial for code translation and these open-source models could also operate locally, assuming the hardware requirements are met.
@@ V-A Limitations
V-B Future Work
In this paper, we assess the quality of the code translation using a binary metric that measures whether the translation was successful, but future work should explore more robust evaluation metrics that penalize more for significant functional errors rather than subtle syntactic errors (e.g., missing imports) that disproportionately affect the binary metric.
Future studies can also expand this work by considering additional languages beyond C++, C, Python, Java, and Go, including older languages such as COBOL, Pascal, and Fortran for legacy code translation. Alternatively, the study can be extended to focus specifically on test generation performance, particularly in creating effective unit and integration tests for production-level code. With the rising popularity of agents, future work can also examine using agentic artificial intelligence to improve code generation quality where the agent automatically corrects syntactic and minor errors that do not affect functional correctness.
Our work demonstrates that the general-purpose GPT-4 8x7B LLM outperformed larger, code-specialized LLMs; however, open-source models still lag behind the reported performance of proprietary models. Future research should investigate the reasons behind the performance gap between general-purpose and code-specialized LLMs, as well as between proprietary and open-source models, to better understand key architectural or training differences.
In addition, experimentation with few-shot learning and variations of CoT prompting (e.g., Structured CoTs [21], COTTON [22]) could also be conducted to determine if intermediate representations would also aid in increasing performance for code translation.
Other future work could be to improve the post-processing of the LLM output along with trying more specific intermediate code representations during code translation. In this work, we post-processed the LLM output to reduce the frequency of noisy non-code lines. However, we only addressed the most common issues using a simple heuristic check, but there could be more specific post-processing to boost the percentage of successful translations. Another lens of future work would be to leverage more specific intermediate code representations. In our prompt formats, we focused on NL and a general AST but there has been progress on self-optimizing ASTs to incorporate specific language nuances [45]. Control flow graphs [46] are another option that could be used as an intermediate code representation.
VI Conclusions